Shown below is an income statement prepared by a poorly trained bookkeeper.

Assume an income tax rate of 30% on all items. Ignore EPS.                

ABC Company

Income Statement

December 31, 20x1

Sales Revenue                                               

$1,192,375

Investment Revenue                                             

9,250

Cost of goods sold                                              

(340,375)

Selling + Administrative                                    

(563,250)

Interest expenses                                                

(13,000)

Loss on Building (Fire Damage)                      

(80,000)

Loss on disposal of business segment            

(32,500)

NI                                                                          

120,750

Prepare a multiple-step income statement in good form. Then answer the questions pertaining to the statement.

XYZ, Inc

Income Statement

December 31, 20x1

Sales Revenue

Cost of Goods Sold

Gross Profit/Margin

Selling and Administrative Expenses

Income from Operations

Other Revenue:

Investment Revenue

Total Other Revenue

Other Expenses:

Interest expense

Loss on building (Fire Damage)

Total Other Expenses

Income from Continuing Operations before Tax

Income Taxes

Income from Continuing Operations

Discontinued Operations:

Loss on Disposal of Business Segment (net of tax of $12,000)

NI

Question 19 options:

What is Total Expenses?

What is Gross Profit/Margin?

What is Total Revenue?

123456789

What is the amount reported for “Discontinued Operations” and does the presentation differ among the single and multi-step format?

What is the amount of “Net Income” and does it differ among the single and multi-step format?

1.

$836,625

2.

$852,000

3.

$1,201,625

4.

It is not displayed.

5.

Loss of $22,750 and yes, it differs.

6.

Loss of $22,750 and no it does not differ.

7.

$120,750 and yes, it differs.

8.

$120,750 and no it does not differ.

9.

Cannot be determined
Answer #1
ABC Company
Income Statement
Particulars Amount Amount
Revenue 1192375
Investment income 9250
Total revenue 1201625
Less: Expenses
Cost of goods sold 340375
Selling & admin expenses 563250
Interest expenses 13000
Loss on building (fire) 80000
Total expenses 996625
Profit from continuing operations 205000
Loss on disposal of business segment -32500
Net income 172500
1. Total expenses= 996625+32500= $109125
2. Gross Margin= Gross Profit/Net sales
Gross profit= Revenue-COGS= 1192375-340375= $852000
Net revenue= 1192375
Gross margin= 852000/1192375= 71.45%
3. Total revenue= Revenue+investment income= $1201625
4. The amount of net income is $ 172500.
It does not differ with preparation of single or multi step income statement
It is only the format of presenting the income statement